Your friend from Netherlands is planning to visit you in the United States. You estimate the cost of her trip at $4,568. What is the cost to her in euros if 1 Euro = $1.4142?
Answer #1

To solve this we need currency exchange rate calculations.

1 Euro = $ 1.4142

Hence,

1 dollar = 1 / 1.4142 Euro

We need to find Euro equivalent to $ 4568

hence,

4568 Dollars = 4568 * (1 / 1.4142) Euro

4568 Dollars = 4568 / 1.4142 Euro

4568 Dollars = 3230.0948 Euro

Hence, cost of trip in Euro = 3230.0948 Euro